Solve the given differential equation using an appropriate substitution. The DE is a bernoulli equation.

t^2 dy/dt + y^2 = ty
1
Expert's answer
2020-07-20T18:13:54-0400

"t^2\frac{dy}{dt}+y^2=ty"

"t^2\frac{dy}{dt}-ty=-y^2"

Dividing equation by "-t^2y^2" :

"-y^{-2}y'+y^{-1}t^{-1}=t^{-2}"

Substituting "z=y^{1-2}=y^{-1}" , "\frac{dz}{dt}=z'=-y^{-2}y'" , therefore: (1)

"z'+\frac z t =\frac 1 {t^2}"

Substituting "z=uv" , "z'=u'v+uv'" : (2)

"u'v+uv'+\frac {uv} t = \frac 1 {t^2}"

"u'v+u(v'+\frac {v} t) = \frac 1 {t^2}" (3)

Setting the part inside () to 0 and separating variables:

"v'+\frac {v} t=0"

"\frac {dv}{dt}=-\frac v t"

"\int \frac {dv} v = - \int \frac {dt} t"

"ln(v) = -ln(t)+ln(k)"

"v=\frac k t" (4)

Putting "v" from (4) into equation (3):

"u'\frac k t = \frac 1 {t^2}"

"\int kdu = \int \frac {dt}t"

"u= \frac 1 k (ln(t)+C)"

Finding z from (2):

"z=uv=\frac 1k (ln(t)+C)\frac kt = \frac{ln(t)+C}{t}"

Finding y from (1):

"y=z^{-1}=\frac t {ln(t)+C}"

Answer: "y=\frac t {ln(t)+C}"